OpenStudy (dls):

P is a variable point on the ellipse x^2/a^2+y^2/b^2= 2 whose foci are F1 and F2.The maximum area of PF1F2 is ?

OpenStudy (dls):

 hey first tell me how do we know is a greater or b ? :/ how do we decide will it be horizontal one or vertical..

OpenStudy (shubhamsrg):

We don't need to know which is greater, just assume any to be greater, form the required equation, everything will be adjusted when values will be given.

OpenStudy (dls):

okay..

OpenStudy (shubhamsrg):

Now, F1 and F2 are fixed, you can find those co-ordinates right ?

OpenStudy (dls):

\[\LARGE b^2=a^2(1-e^2)\] \[\LARGE e=\sqrt{\frac{a^2-b^2}{a^2}}\] and distance between focus=2ae F1=(ae,0) F=(-ae,0)

OpenStudy (dls):

\[\LARGE AREA=\frac{1}{2} 2ae \times y\]

OpenStudy (shubhamsrg):



OpenStudy (shubhamsrg):

Yep, you're on the right track,

OpenStudy (dls):

and y I suppose is \[\LARGE y=b \sqrt{\frac{x^2}{a^2}-2}\]

OpenStudy (shubhamsrg):

y is the y co-ordinate of P right?

OpenStudy (shubhamsrg):

yes y also equals that, but no need of that.

OpenStudy (dls):

yes

OpenStudy (dls):

kaise nikalenge :o

OpenStudy (shubhamsrg):

So, area will be max when y is max.

OpenStudy (shubhamsrg):

y will be max when P will be on the y-axis.

OpenStudy (shubhamsrg):

As easy as that.

OpenStudy (dls):

\[\LARGE A=a \times\sqrt{\frac{a^2-b^2}{a^2}} \times b \sqrt{\frac{x^2}{a^2}-2}\]

OpenStudy (dls):

\[\LARGE A=\sqrt{2}b \sqrt{a^2-b^2}\]

OpenStudy (dls):

Let \[\LARGE P=(\sqrt2acos \phi ,\sqrt2b \sin \phi)\] F1 and F2=(+ - sqrt ae ,0)